OpenStudy (anonymous):

Carrie wanted to make pudding for dessert for her family. However, she couldnt find a recipe that would make 4 servings. She decided to use a recipe for 8 servings and use only 1/2 as much of each ingredient. The recipe for 8 called for 1 1/4 cups of raisins. How many cups of raisins will she need for 4 people? A. 6/8 B. 3/4 C. 2/3 D. 5/8

OpenStudy (anonymous):

@perl @Awolflover1 @pooja195 @zepdrix

OpenStudy (anonymous):

@GirlgoyleH.

OpenStudy (perl):

we can make a proportion

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

You will need 5/8 cups of raisins

OpenStudy (anonymous):

I am trying this on paper and I keep messing up

OpenStudy (anonymous):

The reason why is because you initially need 5/4, but you are trying to make food for only 4 people so that is half of 8 so just simply divide 5/4 by 2 and you get 5/8.

OpenStudy (perl):

$$ \LARGE \frac{1 \frac14~ cup ~raisins}{8~ servings} = \frac{x ~cup~raisins}{4 ~ servings} $$

OpenStudy (perl):

if we use arjund's idea, we have

OpenStudy (anonymous):

okay everytime I try this on paper I am getting thus 1 1/4 + 1/8 = 5/4 x 8/1 = 40/4

OpenStudy (perl):

$$ \\\LARGE \text{since }1 \frac14 = \frac 54 \\ \LARGE \frac{\frac54~ cup ~raisins}{8~ servings} = \frac{\frac{\frac{5}{4}}{\color{red}2} ~cup~raisins}{\frac{8}{\color{red} 2} ~ servings} $$

OpenStudy (anonymous):

okay so it 5/8

OpenStudy (perl):

also you can solve it using a proportion and cross multiplying

OpenStudy (anonymous):

okay

OpenStudy (perl):

$$ \LARGE{\ \frac{\frac54~ cup ~raisins}{8~ servings} = \frac{x ~cup~raisins}{4 ~ servings} \\~\\ \iff \\ \\ 8\cdot x = 4 \cdot \frac 54 \\ \iff \\ 8x = \not 4 \cdot \frac {5}{\not4 } \\ \iff \\ 8x = 5 \\ \iff \\ x = \frac58 } $$

OpenStudy (anonymous):

My idea is made from reasoning, perl's is off math. However, both work. it's up to you on what suite you the best.

OpenStudy (anonymous):

okay thanks so much

OpenStudy (anonymous):

No problem, anything else?

OpenStudy (anonymous):

yes but i will be right back got to see what my mom want

OpenStudy (anonymous):

ok make new post and tag me

OpenStudy (anonymous):

okay